OpenStudy (anonymous):

Show that the equation is not an identity by finding a value of x for which both sides are defined but not equal. cos x - cos x sinx = cos^3 x A. π B. 0 C. (pi)/4 D. (pi)/2

OpenStudy (cwrw238):

i think its x = pi/4 lets check cos x - cosx sinx = 1/sqrt 2 - 1/2 = 0.207 cos^3 x = (1/sqrt2)^3 = 1 / 2^(3/2) = 0.3535 thats the one C

OpenStudy (anonymous):

thnx!